f A and B are any two events of a sample space S, then P.A/ D P.A and B/ 􀀀 P.B/.


1
Expert's answer
2021-06-21T14:39:57-0400

a)

"P(A|B)=\dfrac{P(A\cap B)}{P(B)}"

"=\dfrac{P(A)+P(B)-P(A\cup B)}{P(B)}"

"=\dfrac{P(A)+P(B)-(1-P(\overline{A}\cap \overline{B}))}{P(B)}"

"=\dfrac{P(A)+P(B)-1+P(\overline{A}\cap \overline{B})}{P(B)}"

"P(\overline{A}\cap \overline{B})\geq0"

Hence


"P(A|B)\geq\dfrac{P(A)+P(B)-1}{P(B)}"

The statement is True.


b)


"P(A\cap B)=P(A)-P(A\cap\overline{B})"

Then


"P(A\cap B)=P(A)-P(\overline{A}\cap\overline{B}), \text{ does not hold}"

The statement is True.


c)


"P(A\cup B)=1-P(\overline{A}\cap \overline{B})"

If "A" and "B" are independent, then


"P(\overline{A}\cap \overline{B})=P(\overline{A})P(\overline{B})"

Hence


"P(A\cup B)=1-P(\overline{A}\cap \overline{B})=1-P(\overline{A})P(\overline{B})"

The statement "P(\overline{A}\cap \overline{B})=P(\overline{A})P(\overline{B})," if "A" and "B" are independent, is True.


d) Unless "A" and "B" are mentioned as independent, "P(\overline{A}\cap \overline{B})" cannot be written as "P(\overline{A})P(\overline{B})."

The statement "P(\overline{A}\cap \overline{B})=P(\overline{A})P(\overline{B})," if "A" and "B" are disjoint, is False.